redhoodie:

A conjecture and a portion of the flowchart proof used to prove the conjecture are shown. Which equation best represents the information that should be in the box labeled 1? 2a + 4 = 20 4a = 10 4a + 4 = 20 2a = 20

Vocaloid:

Picture of the flowchart?

redhoodie:

1 attachment
Vocaloid:

the previous step is XY = YZ notice how they also give you XY = 4a + 4 and YZ = 20 according to the substitution property of equality, you can substitute 4a + 4 for XY and 20 for YZ into XY = YZ perform that operation, and you'll have your answer

redhoodie:

thank you very much

redhoodie:

I understand exactly what you mean my question is what operation are you talking about?

Vocaloid:

substitute 4a + 4 for XY and 20 for YZ into XY = YZ

redhoodie:

I see thank you for the clarification so basically you are trying to find out what equation best its for the substitution property of equality. Thank you